# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ Abstract.
Let $G$ be a finite group and let ${\mathrm{cd}}(G)$ be the set of all complex irreducible character degrees of $G.$ In this paper, we show that if ${\mathrm{cd}}(G)={\mathrm{cd}}(H),$ where $H$ is a finite simple exceptional group of Lie type, then $G\cong H\times A,$ where $A$ is an abelian group. This completes the verification of Huppert’s Conjecture for all finite simple exceptional groups of Lie type.
@@ 1. Introduction
Let $G$ be a finite group. Denote by ${\mathrm{Irr}}(G)$ the set of all complex irreducible characters of $G$. Let ${\mathrm{cd}}(G)$ be the set of character degrees of $G$, that is, ${\mathrm{cd}}(G)=\{\chi(1):\chi\in{\mathrm{Irr}}(G)\}.$ One of the main questions in the character theory of finite groups is to determine which properties of a finite group can be determined by its character table. Note that the character degree set of a group can be read off from the character table. Thus it is interesting to ask how much the character degree set of a group knows about the group itself.
Apparently, we do not expect much since the character degree set of a finite group $G$ and the character degree set of a direct product $G\times A$, where $A$ is an abelian group, coincide. In fact, G. Navarro [21] showed that the character degree set alone cannot determine the solvability of the group. He constructed a finite perfect group $H$ and a finite solvable group $G$ such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Furthermore, Navarro and Rizo [23] found a finite perfect group and a finite nilpotent group with the same character degree set.
On the other hand, B. Huppert proposed in [9] that all finite nonabelian simple groups are essentially determined by the set of their character degrees, that is, if $H$ is a finite nonabelian simple group and $G$ is a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$, then $G\cong H\times A$ for some finite abelian group $A$. The main purpose of this paper is to prove the following theorem.
Theorem.
@@
Since the conjecture has been verified for all finite simple exceptional groups of Lie type of Lie rank at most 2, (see [9, 25, 26, 28, 29, 30]), we only need to verify the conjecture for the following finite simple groups: $\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q)$, where $q$ is a prime power.
@@
We should mention that Huppert’s Conjecture has been verified for all sporadic simple groups (see [27] and the references therein), the alternating groups [1] and many finite simple classical groups of small Lie rank (see [11]). The conjecture remains widely open for finite simple classical groups of large rank.
@@
In [12], we extend the conjecture to include certain quasi-simple groups. Recently, an analog of Hupper’s Conjecture was proposed by G. Qian in [14, Problem 20.79] using codegrees. Recall that the codegree of a character $\chi\in{\mathrm{Irr}}(G)$ is defined to be the ratio $|G:{\mathrm{ker}}(\chi)|/\chi(1)$, where ${\mathrm{ker}}(\chi)$ denotes the kernel of $\chi$. Let $\textrm{cod}(G)$ be the set of all codegrees of $G$. It is conjectured by G. Qian that if $G$ is a finite group and $H$ is a nonabelian simple group such that $\textrm{cod}(G)=\textrm{cod}(H)$, then $G\cong H.$ This conjecture has been verified for alternating groups, all sporadic simple groups, all finite simple exceptional groups of Lie type and finite simple projective special linear groups (see [24] and the references therein).
@@
We now describe our approach to the proof of Huppert’s Conjecture. In [9], Huppert proposed a $5$-step strategy to verify his conjecture and this has been used successfully to verify the conjecture for many finite simple groups. In [1], we modified his method to verify the conjecture for alternating groups. We will make use of this method again together with the data from [16] to prove the main theorem.
@@
Suppose that $G$ is a finite group and $H$ is a finite nonabelian simple group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. To verify Huppert’s Conjecture for the nonabelian simple group $H$, we need to prove the following.
Step 1: Show that $G^{\prime}=G^{\prime\prime}$;
Step 2: If $L/M$ is any nonabelian chief factor of $G$, then $L/M\cong H;$
Step 3: If $L$ is a finite perfect group and $M$ is a minimal normal elementary abelian subgroup of $L$ such that $L/M\cong H$, then some degree of $L$ divides no degree of $H$.
Step 4: If $T$ is any finite group with $H\unlhd T\leq\mathop{\rm Aut}\nolimits(H)$ and $T\neq H$, then ${\mathrm{cd}}(T)\nsubseteq{\mathrm{cd}}(H)$.
@@
To verify Steps $1$ and $2$, we will follow the argument developed by Huppert. For Step 3, we first see that either ${\mathbf{C}}_{L}(M)=M$ or ${\mathbf{C}}_{L}(M)=L$. For the first possibility, we use a result due to Guralnick and Tiep [8] on the non-coprime $k(GV)$ problem. If the latter holds, then $L$ is a quasi-simple group. This and Step 4 can now be dealt with using the data from [16].
@@
The paper is organized as follows. We collect some results needed to verify Steps 1 and 3 in Section 2. In Section 3, we present all the information about the character degrees of the finite simple exceptional groups of Lie type that we will need for the proof of the main theorem. We will verify Steps 1–4 in Sections 4–7, respectively.
@@
If $n$ is an integer then we denote by $\pi(n)$ the set of all prime divisors of $n.$ If $G$ is a group, we write $\pi(G)$ to denote the set of all prime divisors of the order of $G.$ Let $\rho(G)=\cup_{\chi\in{\mathrm{Irr}}(G)}\pi(\chi(1))$ be the set of all primes which divide some irreducible character degrees of $G.$ The largest character degree of $G$ will be denoted by $b(G).$ For a group $G,$ let $k(G)$ be the number of conjugacy classes of $G.$ If $N\unlhd G$ and $\theta\in{\mathrm{Irr}}(N),$ then the inertia group of $\theta$ in $G$ is denoted by $I_{G}(\theta).$ The set of all irreducible constituents of $\theta^{G}$ is denoted by ${\mathrm{Irr}}(G|\theta).$ Finally, we denote by $\Phi_{n}=\Phi_{n}(q),$ the cyclotomic polynomial in variable $q.$ Other notation is standard.
@@ 2. Preliminaries
We collect in this section various results that are needed for the proofs of the main theorem. We start with the following lemma which will be used to verify Steps $2$ and $3$
@@
If $S$ is a nonabelian simple group, then there exists a nontrivial irreducible character $\theta$ of $S$ that extends to ${\mathop{\rm Aut}\nolimits}(S).$ Moreover the following holds:
$(i)$ If $S$ is an alternating group of degree at least $7,$ then $S$ has two characters of consecutive degrees $n(n-3)/2$ and $(n-1)(n-2)/2$ that both extend to ${\mathop{\rm Aut}\nolimits}(S).$
$(ii)$ If $S$ is a sporadic simple group or the Tits group, then $S$ has two nontrivial irreducible characters of coprime degrees which both extend to ${\mathop{\rm Aut}\nolimits}(S).$
@@
This is essentially [2, Theorems 2,3,4]. However for sporadic simple groups and the Tits group, we will choose different pairs of coprime degrees. These two degrees are listed in Table 1. ∎
We also need the following number theoretic result. It will be used to show that certain character degree of the simple exceptional groups of Lie type cannot be written as a proper nontrivial power.
@@
The equation in the lemma can be rewritten as $(x^{3}-1)/(x-1)=y^{m}$. This is a special case of the Nagell-Ljunggren equation $(x^{n}-1)/(x-1)=y^{m}.$ Nagell and Ljunggren show that (see [3]) if $n=3$, then the only solution of the equation with $x,y,m\geq 2$ is $(x,y)=(18,7)$. However, $x=18$ is not a prime power. ∎
The next result is crucial to our proof of Step $3$ Unfortunately, we are unable to find a simple proof of this fact. The proof given below is an easy consequence of a deep result due to R. Guralnick and P.H. Tiep [8] concerning the non-coprime $k(GV)$-problem. For a finite group $G,$ we let $b(G)$ be the largest degree of $G.$ It follows from character theory that $|G|=\sum_{\chi\in{\mathrm{Irr}}(G)}\chi(1)^{2}$ and $k(G)=|{\mathrm{Irr}}(G)|,$ where $k(G)$ is the number of conjugacy classes of $G.$ Thus $|G|\leq k(G)b(G)^{2}$ and $b(G)^{2}\leq|G|.$
@@
Let $H\cong L/M$ be a finite simple exceptional group of Lie type of Lie rank at least $4$ By way of contradiction, assume that $b(L)\leq b(L/M)=b(H).$ Form the semidirect product $HM$ with $H$ acting on $M$ as inside $L.$ It follows from [8, Lemma 2.4] that $k(L)\leq k(HM).$ We now have that $|L|=\sum_{\chi\in{\mathrm{Irr}}(L)}\chi(1)^{2}\leq k(L)b(L)^{2}\leq k(HM)b(H)^{2}.$ As $H$ is a finite simple exceptional group of Lie type of Lie rank at least $4,$ and $H$ acts irreducibly on $M,$ by applying [8, Theorems 1.4, 1.5], we have $k(HM)\leq|M|/2.$ Furthermore $|L|=|H|\cdot|M|$ and so the above inequality becomes $|L|=|H|\cdot|M|\leq|M|b(H)^{2}/2.$ After simplifying, we obtain that $|H|\leq b(H)^{2}/2<b(H)^{2},$ which is impossible. Therefore, $b(L)>b(L/M)$ as wanted. ∎
To verify Step 3 for classical groups, we will need the following result.
Problem 2.4.
@@
It is even possible that $b(G)>b(G/M)$ in the above situation. To finish Step $3$, we still need to show that if $L$ is a quasisimple group with $L/Z(L)$ a finite simple classical group, then some character degree of $L$ divides no degree of $L/Z(L)$.
@@
([15, Theorems  5.2.14, 5.2.15]). Let $q$ and $n$ be integers with $q\geq 2,$ and $n\geq 3.$ Suppose that $(q,n)\neq(2,6).$ Then $q^{n}-1$ has a prime divisor $\ell$ such that $\ell\equiv 1$ (mod $n$) and $\ell$ does not divide $q^{k}-1$ for any $k<n.$ Moreover if $\ell\mid q^{k}-1$ for some integer $k,$ then $n\mid k.$
Such a prime $\ell$ is called a primitive prime divisor (ppd for short) and is denoted by $\ell_{n}(q).$ When $n$ is odd and $(q,n)\neq(2,3)$ then there is a primitive prime divisor of $q^{2n}-1$ and we denote this primitive prime divisor by $l_{-n}(q).$ In terms of the cyclotomic polynomials, as $q^{n}-1=\prod_{k\mid n}\Phi_{k},$ we obtain that if $\ell$ is a primitive prime divisor of $q^{n}-1,$ then $\ell\mid\Phi_{n}$ but $\ell\nmid\Phi_{k}$ for any $k<n.$ In this situation, we also say that $\ell$ is a primitive prime divisor of $\Phi_{n}.$
@@ 3. Character degrees of simple exceptional groups of Lie type
In this section, we collect all the results concerning the character degrees of exceptional groups of Lie type that we need for the proof of Theorem Theorem. We first consider the following set up. Let $\mathcal{G}$ be a simple simply connected algebraic group defined over the algebraic closure of a finite field ${\mathbb{F}}_{p}$ of characteristic $p$ and let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a Frobenius morphism such that the fixed point group $L=\mathcal{G}^{F}$ is finite and $H\cong L/Z(L)$ is a finite simple groups. Let $\mathcal{G}^{*}$ be a group dual to $\mathcal{G}$ with a dual Frobenius map $F^{*}$ and let $H^{*}=({\mathcal{G}^{*}})^{F^{*}}.$ We also use the notation $H_{sc}$ for $L$ and $H_{ad}$ for $H^{*}.$
According to Lusztig, there is a natural partition
@@
of ${\mathrm{Irr}}(L),$ where the union runs over $H^{*}$-conjugacy classes of semisimple elements. The irreducible characters of $L$ in $\mathcal{E}(L,1)$ are call unipotent characters of $L.$ The unipotent character degrees of simple exceptional groups of Lie type are given explicitly in [4, 13.9]. Moreover for each $s,$ there is a bijection map
@@
such that for any $\chi\in\mathcal{E}(L,s),$ the degree of $\chi$ is given by
@@
By results of Lusztig, we know that all the unipotent characters of $L$ are trivial on the center of $L,$ therefore they are exactly all the unipotent characters of $H\cong L/Z(L).$ Moreover, by result of G. Malle in [17], all unipotent characters of $H$ extend to $\mathop{\rm Aut}\nolimits(H),$ apart from some exceptions (see [17, Theorem 2.5]). This fact will be used frequently in the proofs of Steps $2$ and $3$ Some results on semisimple characters of $L$ can be found in [6, Lemma 2.5]. In Table 2, we list the $p$-parts of the degrees of some unipotent characters of finite simple groups of Lie type in characteristic $p$.
@@
All the results in this section are based on the data given in [16]. From this, we can extract all the information about the character degrees of both $L$ and $H^{*},$ the simply connected and adjoint versions of the simple exceptional group of Lie type $H.$ Note that if $d:=|Z(L)|=|H^{*}:H|$ is trivial, then all these versions of $H$ are isomorphic and we will obtain the character degrees of the simple group $H.$ Otherwise, we only know that ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L)$ and every character degree of $H$ must divide some degree of $H^{*}.$

### v3

@@ Abstract.
Let $G$ be a finite group and let ${\mathrm{cd}}(G)$ be the set of its complex irreducible character degrees. We show that if ${\mathrm{cd}}(G)={\mathrm{cd}}(H),$ where $H$ is a finite simple exceptional group of Lie type, then $G\cong H\times A,$ where $A$ is an abelian group. This completes the verification of Huppert’s Conjecture for all finite simple exceptional groups of Lie type.
@@ 1. Introduction
Let $G$ be a finite group and let ${\mathrm{Irr}}(G)$ be the set of its complex irreducible characters. By a ‘character degree’ we mean the degree of a complex irreducible character. Let ${\mathrm{cd}}(G)$ be the set of all character degrees of $G$. One of the main questions in the character theory of finite groups is to determine which properties of a finite group can be determined by its character table. Note that the character degree set of a group can be read off from the character table. Thus it is interesting to ask how much the character degree set of a group knows about the group itself.
We cannot expect too much since the character degree set of a finite group $G$ and the character degree set of a direct product $G\times A$, where $A$ is an abelian group, coincide. Navarro [25] showed that the character degree set alone cannot determine the solvability of the group. He constructed a finite perfect group $H$ and a finite solvable group $G$ such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Furthermore, Navarro and Rizo [27] found a finite perfect group and a finite nilpotent group with the same character degree set.
On the other hand, Huppert proposed in [12] that all finite nonabelian simple groups are essentially determined by the set of their character degrees: if $H$ is a finite nonabelian simple group and $G$ is a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$, then $G\cong H\times A$ for some finite abelian group $A$. The main purpose of this paper is to prove the following theorem.
Theorem 1.1.
@@
Huppert’s Conjecture has been verified for various finite simple exceptional groups of Lie type (see [12, 29, 30, 32, 34]). To complete the verification of this conjecture for finite simple exceptional groups of Lie type, we verify the conjecture for the following finite simple groups: $\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q)$ where $q$ is a prime power.
@@
Huppert’s Conjecture has been verified for all sporadic simple groups (see [31] and the references therein), the alternating groups [1] and many finite simple classical groups of small Lie rank (see [14]). The conjecture remains open for finite simple classical groups of large rank.
@@
In [15], we extend the conjecture to include certain quasi-simple groups. Recently, an analog of Huppert’s Conjecture was proposed by Qian in [17, Problem 20.79] using codegrees. Recall that the codegree of a character $\chi\in{\mathrm{Irr}}(G)$ is the ratio $|G:{\mathrm{ker}}(\chi)|/\chi(1)$, where ${\mathrm{ker}}(\chi)$ denotes the kernel of $\chi$. Let $\textrm{cod}(G)$ be the set of all codegrees of $G$. Qian conjectured that if $G$ is a finite group and $H$ is a finite nonabelian simple group such that $\textrm{cod}(G)=\textrm{cod}(H)$, then $G\cong H.$ This conjecture has been verified for alternating groups, all sporadic simple groups, all finite simple exceptional groups of Lie type and finite simple projective special linear groups (see [28] and the references therein).
@@
We now describe our approach to the proof of Huppert’s Conjecture. In [12], Huppert proposed a $5$-step strategy to verify his conjecture and this has been used successfully to verify the conjecture for many finite simple groups. In [1], we modified his method to verify the conjecture for alternating groups. We use this method and the data from [20] to prove the main theorem.
@@
Suppose that $G$ is a finite group and $H$ is a finite nonabelian simple group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. To verify Huppert’s Conjecture for $H$, we need to prove the following.
Step 1: Show that $G^{\prime}=G^{\prime\prime}$.
Step 2: If $G^{\prime}/M$ is a chief factor of $G$, then $G^{\prime}/M\cong H$.
Step 3: Show that $M=1$.
Step 4: Show that $G\cong G^{\prime}\times C_{G}(G^{\prime})$.
@@
To verify Steps $1$ and $2$, we follow the argument developed by Huppert. For Step 3, we can reduce to the case when $L$ is a quotient of $G^{\prime}$ and $N$ is a minimal normal elementary abelian subgroup of $L$ such that $L/N\cong H$ and either ${\mathbf{C}}_{L}(N)=N$ or ${\mathbf{C}}_{L}(N)=L$ (see the proof of Proposition 6.1). For the first possibility, we use a result due to Guralnick and Tiep [11] on the non-coprime $k(GV)$-problem to obtain a contradiction. If the latter holds, then $L$ is a quasi-simple group. This and Step 4 can now be dealt with using the data from [20].
@@
The paper is organized as follows. We collect some results needed to verify Steps $1,2$ and $3$ in Section 2. In Section 3, we present all the information about the character degrees of the finite simple exceptional groups of Lie type that we need for the proof of the main theorem. We verify Steps 1–4 in Sections 47, respectively.
@@
If $n$ is an integer, then we denote by $\pi(n)$ the set of all prime divisors of $n.$ If $G$ is a group, then we write $\pi(G)$ to denote the set of all prime divisors of the order of $G.$ If $p$ is a prime and $n>1$ is an integer, then we write $n_{p}$ for the largest power of $p$ that divides $n$. Let $\rho(G)=\cup_{\chi\in{\mathrm{Irr}}(G)}\pi(\chi(1))$ be the set of all primes which divide some irreducible character degrees of $G.$ The largest character degree of $G$ is denoted by $b(G).$ For a group $G,$ let $k(G)$ be the number of conjugacy classes of $G.$ If $N\unlhd G$ and $\theta\in{\mathrm{Irr}}(N),$ then the inertia group of $\theta$ in $G$ is denoted by $I_{G}(\theta).$ The set of all irreducible constituents of $\theta^{G}$ is denoted by ${\mathrm{Irr}}(G|\theta).$ Finally, we denote by $\Phi_{n}=\Phi_{n}(q)$ the cyclotomic polynomial in variable $q.$ Other notation is standard.
@@ 2. Preliminaries
We collect various results that are needed for the proofs of the main theorem. We start with the following lemmas which will be used to verify Steps $2$ and $3$
@@
If $S$ is a nonabelian simple group, then there exists a nontrivial irreducible character $\theta$ of $S$ that extends to ${\mathop{\rm Aut}\nolimits}(S).$ Moreover the following hold:
$(i)$ If $S$ is an alternating group of degree at least $7,$ then $S$ has two characters of consecutive degrees $n(n-3)/2$ and $(n-1)(n-2)/2$ that extend to ${\mathop{\rm Aut}\nolimits}(S).$
$(ii)$ If $S$ is a sporadic simple group or the Tits group, then $S$ has two nontrivial irreducible characters which extend to ${\mathop{\rm Aut}\nolimits}(S)$ and have coprime degrees. These characters are listed in Table 1,
@@
This is essentially [2, Theorems $2,3,4$]. However for sporadic simple groups and the Tits group, we choose different pairs of coprime degrees from those in [2]. ∎
Lemma 2.1 will be used together with the following extension result.
Lemma 2.2.
Let $N$ be a minimal normal subgroup of a finite group $G$ such that $N\cong S^{k},$ where $S$ is a nonabelian simple group. If $\theta\in{\mathrm{Irr}}(S)$ extends to ${\rm{Aut}}(S),$ then $\theta^{k}\in{\mathrm{Irr}}(N)$ extends to $G.$
Proof.
This is Lemma $5$ in [2]. ∎
We also need the following number theoretic result. It is used to show that a certain character degree of the simple exceptional groups of Lie type cannot be written as a proper nontrivial power.
@@
The equation in the lemma can be rewritten as $(x^{3}-1)/(x-1)=y^{m}$. This is a special case of the Nagell-Ljunggren equation $(x^{n}-1)/(x-1)=y^{m}.$ Nagell and Ljunggren show (see [3]) that if $n=3$, then the only solution of the equation with $x,y,m\geq 2$ is $(x,y)=(18,7)$. However, $x=18$ is not a prime power. ∎
We next list the irreducible characters of prime power degrees of the finite nonabelian simple exceptional groups of Lie type. This is part of [23, Theorem 1.1].
Lemma 2.4.
Let $G$ be one of the finite simple exceptional groups of Lie type listed in (1) and let $\chi$ be a faithful irreducible character of $G$. Suppose that $\chi(1)=r^{d}$, where $r$ is a prime. Then $G$ is of characteristic $r$ and $\chi$ is the Steinberg character of $G$, so $\chi(1)=|G|_{r}$.
The next result is crucial to our proof of Step $3$ Unfortunately, we are unable to find a simple proof of this fact. The proof given below is an easy consequence of a deep result due to Guralnick and Tiep [11] concerning the non-coprime $k(GV)$-problem. For a finite group $G,$ let $b(G)$ be the largest character degree of $G.$ It follows from character theory that $|G|=\sum_{\chi\in{\mathrm{Irr}}(G)}\chi(1)^{2}$ and $k(G)=|{\mathrm{Irr}}(G)|,$ where $k(G)$ is the number of conjugacy classes of $G.$ Thus $|G|\leq k(G)b(G)^{2}$ and $b(G)^{2}\leq|G|.$
@@
Let $H\cong L/M$ be a finite simple exceptional group of Lie type of Lie rank at least $4$ By way of contradiction, assume that $b(L)\leq b(L/M)=b(H).$ Since $M$ is an abelian normal subgroup of $L$, the conjugation action of $L$ on $M$ induces an action of $H$ on $M$. Using this action, we form the semidirect product $HM$. By [11, Lemma 2.4], $k(L)\leq k(HM).$ Now
$|L|=\sum_{\chi\in{\mathrm{Irr}}(L)}\chi(1)^{2}\leq k(L)b(L)^{2}\leq k(HM)b(H)^{2}.$
As $H$ is a finite simple exceptional group of Lie type of Lie rank at least $4,$ and $H$ acts irreducibly on $M$. By [11, Theorems 1.4, 1.5], $k(HM)\leq|M|/2.$ Moreover, $|L|=|H|\cdot|M|$ and so the above inequality becomes
$|L|=|H|\cdot|M|\leq|M|b(H)^{2}/2.$
Hence $|H|\leq b(H)^{2}/2<b(H)^{2},$ a contradiction. Thus $b(L)>b(L/M)$ as wanted. ∎
To verify Step 3 for finite simple classical groups, we need the following result.
Problem 2.6.
@@
It is possible that $b(G)>b(G/M)$ in the above situation. To finish Step $3$, we need to show that if $L$ is a finite quasi-simple group with $L/Z(L)$ a finite simple classical group, then some character degree of $L$ divides no character degree of $L/Z(L)$.
@@
([18, Theorems  5.2.14, 5.2.15]). Let $q$ and $n$ be integers with $q\geq 2,$ and $n\geq 3.$ Suppose that $(q,n)\neq(2,6).$ Then $q^{n}-1$ has a prime divisor $\ell$ such that $\ell\equiv 1$ (mod $n$) and $\ell$ does not divide $q^{k}-1$ for any $k<n.$ Moreover if $\ell\mid q^{k}-1$ for some integer $k,$ then $n\mid k.$
Such a prime $\ell$ is a primitive prime divisor (ppd for short) and is denoted by $\ell_{n}(q).$ If $n$ is odd and $(q,n)\neq(2,3)$ then $q^{2n}-1$ has a primitive prime divisor denoted by $\ell_{-n}(q).$ In terms of the cyclotomic polynomials, as $q^{n}-1=\prod_{k\mid n}\Phi_{k},$ if $\ell$ is a primitive prime divisor of $q^{n}-1,$ then $\ell\mid\Phi_{n}$ but $\ell\nmid\Phi_{k}$ for any $k<n.$ In this situation, we also say that $\ell$ is a primitive prime divisor of $\Phi_{n}.$
@@ 3. Character degrees of simple exceptional groups of Lie type
We now collect the results concerning the character degrees of exceptional groups of Lie type that we need for the proof of Theorem 1.1. We first consider the following setup (see, for example, [22]). Let $\mathcal{G}$ be a simple simply connected algebraic group defined over the algebraic closure of a finite field ${\mathbb{F}}_{p}$ of characteristic $p$ and let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a Steinberg endomorphism such that the fixed point group $L=\mathcal{G}^{F}$ is finite and $H\cong L/Z(L)$ is a finite simple group. Let $\mathcal{G}^{*}$ be a group dual to $\mathcal{G}$ with a dual Frobenius map $F^{*}$ and let $H^{*}=({\mathcal{G}^{*}})^{F^{*}}.$ We use the notation $H_{sc}$ for $L$ and $H_{ad}$ for $H^{*}.$
According to Deligne - Lusztig theory (see [4, 7, 9]), there is a natural partition
@@
of ${\mathrm{Irr}}(L),$ where the union runs over $H^{*}$-conjugacy classes of semisimple elements. The irreducible characters of $L$ in $\mathcal{E}(L,1)$ are unipotent characters of $L.$ The unipotent character degrees of simple exceptional groups of Lie type are given explicitly in [4, 13.9]. Moreover for each $s,$ there is a bijection map
@@
such that for each $\chi\in\mathcal{E}(L,s),$ the degree of $\chi$ is given by
@@
By results of Lusztig (see [21]), we know that all the unipotent characters of $L$ are trivial on the center of $L$; therefore, they are exactly all the unipotent characters of $H\cong L/Z(L).$ Moreover, unipotent characters of $H$ extend to $\mathop{\rm Aut}\nolimits(H),$ apart from some exceptions (see [21, Theorem 2.5]). This fact will be used frequently in the proofs of Steps $2$ and $3$ Some results on semisimple characters of $L$ can be found in [8, Lemma 2.5]. In Table 2, we list the $p$-parts of the degrees of some unipotent characters of finite simple groups of Lie type in characteristic $p$. Note that we follow the ATLAS notation [6] for finite simple groups of Lie type.
@@
The results in this section are based on the data given in [20]. From this, we can extract all the information about the character degrees of both $L$ and $H^{*},$ the simply connected and adjoint versions of the simple exceptional group of Lie type $H.$ Note that if $d:=|Z(L)|=|H^{*}:H|$ is trivial, then all these versions of $H$ are isomorphic and we obtain the character degrees of the simple group $H.$ Otherwise, we only know that ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L)$ and every character degree of $H$ divides some degree of $H^{*}.$